# Global calibration of large-scale photonic integrated circuits

### v2

@@ Abstract
The advancing maturity of photonic integrated circuit (PIC) fabrication technology enables the high integration of an increasing number of optical components onto a single chip. With the incremental circuit complexity, the calibration of active phase shifters in a large-scale PIC becomes a crucially important issue. The traditional one-by-one calibration techniques encounter significant hurdles with the propagation of calibration errors, and achieving the decoupling of all phase shifters for independent calibration is not straightforward. To address this issue, we propose a global calibration approach for large-scale PIC. Our method utilizes a custom network to simultaneously learn the nonlinear phase-current relations for all thermo-optic phase shifters on the PIC by minimizing the negative likelihood of the measurement datasets. Moreover, the reflectivities of all static beamsplitter components can also be synchronizedly extracted using this calibration method. As an example, a quantum walk PIC with a circuit depth of 12 is calibrated, and a programmable discrete-time quantum walk is experimentally demonstrated. These results will greatly benefit the applications of large-scale PICs in photonic quantum information processing.
@@ I Introduction
The traditional calibration method requires the decoupling of all phase shifters followed by individual calibration for each one [19, 20, 21, 22, 23]. However, this approach isn’t ideal for large-scale PICs due to two main challenges. Firstly, while the traditional method can be used to calibrate the standard triangular “Rech” PICs [24, 25], rectangular “Clements” PICs [26, 27, 28, 29], and PICs with cascaded phase shifters [30, 31] by following a specific decoupling order, it is difficult to find an effective and direct way to decouple all phase shifters in the intricate mesh structures. Secondly, deviations in the calibration values of phase shifters in preceding levels may also spread errors in the calibration of subsequent phase shifters. As a result, calibration errors accumulate progressively with the increasing number of phase shifters, eventually becoming significant and non-negligible.
To address these issues, we present an efficient calibration method developed to globally calibrate all optical components on the PICs, thereby mitigating the accumulation of errors inherent in the traditional one-by-one calibration approach. The global calibration method utilizes a custom network with simulation layers that creates a virtual duplicate of the physical PICs, modelling their practical non-unitary processes involving optical loss. Using the gradient-descent approach, we can determine the optimal values of parameters that need to be calibrated, ensuring that the outputs of the network closely resemble the distributions of measurement data. As an example, on a large-scale quantum walk (QW) PIC with a circuit depth of 12, we find that the trained custom network enables the simultaneous extraction of not only the nonlinear phase-current relations of all the thermo-optic phase shifters but also the reflectivities of all the beamsplitters. Moreover, we further experimentally demonstrate the silicon PIC calibration and implement the on-chip QW [32, 33, 34, 35, 36, 37, 38, 39, 40].
This article is organized as follows: In Sec. II, we will introduce our custom-architecture global calibration method, provide a concise overview of the QW PIC, and evaluate the performance of our calibration method using synthetic datasets featuring Gaussian noise. In Sec. III, we will showcase the calibration results of a silicon PIC utilizing noisy experimental datasets and discuss the performance regarding the experimental implementation of QW. Finally, Sec. IV presents the summary of our work and engages in a discussion.
@@ II.1 Global calibration method
where $\sin^{2}\theta$ denotes the reflectivity of the beamsplitter. Considering the inevitable optical loss in a practical chip, the transfer matrix for a lossy BS becomes non-unitary and is given by [41]:
@@
where $\alpha$ denotes the transfer efficiency of a single BS, and $1-\alpha$ represents its optical loss. The phase shifter provides a relative phase difference $\phi$ between the two output paths of the beamsplitter, and its matrix representation is $\hat{U}_{\text{PS}}=\text{diag}\{1,e^{i\phi}\}$. For the widely-used thermo-optic phase shifter, the relative phase $\phi$ can be tuned by the current $I$ applied to the phase shifter, according to the following formula [13]:
@@
where $a$ and $b$ are hardware parameters to be determined. Then, the non-unitary process $\hat{U}(\bm{a},\bm{b},\bm{\theta},\bm{\alpha})$ of PICs is the matrix product of its individual components. Moreover, considering the mode-field diameter mismatch between the output ports of the PICs and the optical fiber arrays, it is necessary to introduce another parameter, $\bm{\eta}$, to indicate the fiber-chip coupling efficiency. Consequently, $\bm{\lambda}=\{\bm{a},\bm{b},\bm{\theta},\bm{\alpha},\bm{\eta}\}$ are all undetermined parameters for a linear photonic circuit.
@@
The learning process necessitates a number of training data, comprising randomly generated current distributions $\{\bm{I}\}_{n}$ applied to all the phase shifters and the target probability distributions $\{\bm{P}_{\text{tar}}\}_{n}$ of the output ports on the chip. The index $n=1,\cdots,N$ represents each training sample, and $N$ is the total number of training datasets. We set the current distribution $\bm{I}$ as inputs, which are then fed to the input layer of our custom network. The custom-network-parameterized non-unitary process $\hat{U}_{\text{ML}}(\bm{a},\bm{b},\bm{\theta},\bm{\alpha})$ can be determined using Eqs. (2) and (3). For a given initial state $\rho_{i}$, the final state is given by $\rho_{f}=\hat{U}_{\text{ML}}\rho_{i}\hat{U}_{\text{ML}}^{\dagger}$, and the probability distribution $\text{Tr}[\left|v\right\rangle\!\!\left\langle v\right|\rho_{f}]$ can be obtained. Considering the relative fiber-coupling efficiency $\bm{\eta}$ between the output ports of the PICs, the final probability distribution is normalized, and then the parameterized network distribution $\bm{P}_{\text{ML}}(\bm{\lambda})$ is generated.
We define the cost function $\mathcal{L}$ as the summation of the L1-norm distance [42] between the target probability distributions in the training datasets $\{\bm{P}{{}_{\text{tar}}}\}_{n}$ and the parameterized network distributions $\{\bm{P_{\text{ML}}(\bm{\lambda})}\}_{n}$: $\mathcal{L}(\bm{\lambda})=\sum_{n}\frac{1}{2}|\bm{P{{}_{\text{ML}}}}(\bm{\lambda})-\bm{P_{\text{tar}}}|$. The aim of training the custom network is to find the optimal parameters $\bm{\lambda}^{\ast}$ by minimizing the cost function such that the parameterized network distributions approximate the target ones. Then, the parameterized non-unitary process $\hat{U}_{\text{ML}}$ together with the relative coupling efficiency $\bm{\eta}^{\ast}$ for the trained network can well reflect the target physical process that generates the distributions $\{\bm{P_{\text{ML}}}(\bm{\lambda}^{\ast})\}_{n}\simeq\{\bm{P}{{}_{\text{tar}}}\}_{n}$.
@@ II.2 Model of on-chip quantum walks
Here, the parameter $\alpha$ determines the loss strength and varies depending on both the position and time step. $\cos^{2}\theta$ and $\sin^{2}\theta$ denote the probabilities of a walker moving backwards and forwards on the 1D line, respectively. Parameters $\beta$ and $\gamma$ introduce the relative phase difference. The coin operator $\hat{C}$ can be position- and time-dependent, a feature necessary for specific tasks in QW-based quantum simulations [35] and quantum algorithms [14].
For the QW PIC, the coin state {$\{\ket{\uparrow},\ket{\downarrow}\}$ in the QW model can be cast on the path state (inset of Fig. 2): the upper path of the beamsplitter corresponds to |$\ket{u}$, while the lower path corresponds to |$\ket{l}$. The localized coin state at each position is described by |$\ket{c}=c_{1}\ket{u}+c_{2}\ket{l}$, which can be interpreted as a superposition state located on the two paths of the corresponding beamsplitter. Each beamsplitter with a splitting ratio of 50:50 and phase shifter work together (black circle in Fig. 2) to realize the coin operator $\hat{C}$ in the sense of path states, corresponding to Eq. (4) with $\theta=\pi/4$ and $\beta=0$. The interwoven mesh architecture of the PIC, composed of the on-chip waveguides (black solid lines), realizes the shift operators $\hat{S}$. After $T$-step non-unitary dynamics of the QW PIC, the size of the position space is $T$, with each position having two paths representing the coin space. Therefore, the total number of output ports of the QW PIC is $2T$.
@@ II.3 Numerical benchmarks
During the iteration process, simultaneously updating all network parameters using the gradient-descent method is susceptible to getting trapped in a local minimum of $\mathcal{L}$. To mitigate this issue and escape local minima, we utilize a non-simultaneous parameter update method, updating only partial parameters at each epoch while keeping the remaining parameters fixed. The relative coupling efficiency $\bm{\eta}$ is configured to be constantly updated. The reflectivity parameter $\bm{\theta}$ and the relative transfer efficiency $\bm{\alpha}$ are frozen at first. We then alternatively train one of the two parameters $\{\bm{a},\bm{b}\}$ while the remaining one is kept frozen and untrainable. Within a single alternating training process, the training for each parameter ceases upon meeting the cut-off condition, where five consecutive decreases in the loss function between adjacent epochs are each less than $10^{-5}$. After eight rounds of alternating training, we proceed to train the last two parameters, $\bm{\theta}$ and $\bm{\alpha}$, while maintaining the fixed values of $\bm{a}$ and $\bm{b}$. In the end, following the completion of the non-simultaneous parameter updates described above, we proceed to simultaneously update all parameters $\bm{\lambda}$ until the cut-off condition is met.
@@ III.1 Experimental setup
In our experiment, the conceptual diagram of the QW PIC depicted in Fig. 2 is implemented through silicon waveguide circuits fabricated using the complementary metal–oxide semiconductor (CMOS)-compatible silicon photonics process [51], as illustrated in Fig. 5. The $6.8\,\text{mm}\times 3.4\,\text{mm}$ chip comprises an array of 66 individually controlled thermo-optic phase shifters with 78 waveguide beamsplitters to form the cascade of the interference network of a 12-step on-chip QW. The controlled phase shifter is implemented by electrically heating a resistor to change the refractive index of the waveguide to deform the optical path length [52]. Then, the phase shifter has a linear phase-power relation, and its phase change $[0,2\pi]$ can be tuned by changing the imposed current according to the nonlinear phase-current relation in Eq. (3). To program all these phase shifters, we have developed a multi-channel electronic control system. The beamsplitter elements are realized using multimode interference couplers. On-chip grating couplers [53] are used for fiber-chip coupling, and the coupling loss is 4.5 dB per facet. A continuous-wave laser (TLX1, Thorlabs) operating at a wavelength of 1550 nm and with a power output of 5 mW is coupled into the QW PIC through a single-mode optical fiber. The input fiber is precisely aligned with the lower path of the first waveguide beamsplitter using a 3-axis motorized positioning stage (KWL06050, Suruga), and the polarization of the input laser is fine-tuned using the 3-paddle fiber polarization controller (FPC030, Thorlabs). We simultaneously measure the light intensity from all output ports by coupling them to an array of optical fibers and subsequently directing them to Germanium-amplified detectors (PDA30B2, Thorlabs). For each output port, we detect the light intensity five times and store the average value. Consequently, the measured intensity distributions are normalized to the experimental probability distribution $\bm{P}_{\text{exp}}$. In the experiment, the QW PIC is temperature-controlled at 22C.
@@ III.2 Experimental results
Herein, the calibration of the QW PIC is implemented and tested by training the custom network directly on actual experimental datasets $\{\bm{P}_{\text{exp}}\}_{n}$. In order to generate the experimental training and test datasets, we randomly generate 1500 sets of current distributions $\{\bm{I}\}_{n}$ ranging from 0 mA to 7 mA imposed on the phase shifters on our silicon PIC. We obtain 1500 sets of experimental probability distributions $\{\bm{P}_{\text{exp}}\}_{n}$ after normalizing the measured light intensity over all output ports. The experimental datasets are divided into two parts according to an 85:15 ratio, where 1275 sets of datasets are used to train the custom network, and the remaining 225 sets of datasets are used to test the trained network. With these experimental training sets, the parameters $\bm{\lambda}=\{\bm{a},\bm{b},\bm{\theta},\bm{\alpha},\bm{\eta}\}$ of the custom network in Fig. 1 are then pre-trained.
The trained network can characterize the phase-current relation, assumed to be of the form $\phi=aI^{2}+b$, the reflectivity parameter $\bm{\theta}$ and the relative transfer efficiency $\bm{\alpha}$ of BSs, as well as the relative output coupling efficiency $\bm{\eta}$ between the output ports. Here, we provide the relative transfer efficiency between all BSs and the relative coupling efficiency between all output ports because we utilize the normalized probability distributions to calibrate the QW PIC. Even so, we only need to measure the absolute transfer efficiency for one BS and the absolute coupling efficiency for one output port. The efficiencies of all the remaining BSs and output ports can then be directly calculated.
@@
Figure 7: Validation of the QW PIC. (a) The measured probability distribution $\bm{P}^{\text{QW}}_{\text{exp}}$ (blue bars) for 12-step Hadamard QWs. The red bars mean the distribution $\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})$ predicted by the trained network. (b) and (c) display histograms of the measured values of infidelity and purity error for 100 output quantum states after 9-step QW dynamics, respectively.
@@
After completing the calibration of the photonic chip using the training datasets, we utilize the experimental test datasets to further verify the accuracy of the calibration. We apply the remaining 225 sets of distributions of random currents, and compare the experimentally measured distributions $\{\bm{P}^{\text{test}}_{\text{exp}}\}_{n}$ with the predicted probability distributions $\{\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})\}_{n}$ of the trained network with the optimal hyperparameters $\bm{\lambda}^{\ast}$. The difference between the two distributions for each test sample $n$ is defined by L1-norm distance: $d(\bm{\lambda}^{\ast})=\frac{1}{2}|\bm{P{{}_{\text{ML}}}}(\bm{\lambda}^{\ast})-\bm{P}^{\text{test}}_{\text{exp}}|$. We observe that the distance values for $85\%$ of the test samples are less than 0.05, as shown in Fig. 6(f). The average distance for all test samples is $\bar{d}=0.033\pm 0.018$, indicating that our calibration results for the phase shifters and beamsplitters are in good agreement with the actual physical process. The slight deviation suggests the presence of additional chip imperfections that are not considered by our custom network, such as the non-dominant thermal crosstalk effect [54]. The error bar of $\bar{d}$ represents one standard deviation.
@@
After verifying the performance on numerous random validation samples using the calibrated photonic chip, we proceed to implement the conventional 1D Hadamard QWs, where the coin operators are fixed as the Hadamard gate and independent of time and position. We inject 1550 nm light into the lower path of the first waveguide beamsplitter followed by setting the phase of the first phase shifter to $\pi/2$. Therefore, we prepare a balanced superposition initial state between two paths with a phase difference of $\pi/2$ through one-step QW dynamics [55]. The phases of the remaining phase shifters are set to 0. We measure the light intensity coupled to the optical fiber arrays, and the experimental probability distribution $\bm{P}^{\text{QW}}_{\text{exp}}$ for a 12-step QW is obtained using the calibrated fiber-coupling efficiency for compensation, as depicted in Fig. 7(a). The resulting L1-norm distance with respect to the QW distribution predicted by the trained network is $d=0.057$, indicating their good agreement.
@@
Beyond the position distributions, we then perform the quantum state tomography for the calibrated QW PIC with inherent high-dimensional structures [56, 57, 58, 59]. To achieve this, in addition to the computational basis |$\ket{v}$, extra measurement bases need to be introduced to fully characterize the output quantum state. We divide the QW PIC into two sections: the first part, with a circuit depth of 9, is used to implement arbitrary 9-step QW dynamics, while the second part, with a circuit depth of 3, serves as a high-dimensional measurement device [60, 61, 62, 63]. For each 9-step QW dynamics, we impose 14 sets of current distributions into the measurement device and measure the output probability distributions. Here, we perform standard maximum-likelihood reconstruction with the estimator taking the form ${\rho_{\text{m}}=TT^{\dagger}/\text{Tr}(TT^{\dagger})}$, where $T$ is a lower triangular matrix [64]. The quantum states after the 9-step QW can be reconstructed by maximizing the likelihood of 14 sets of experimental distributions. We arbitrarily generate 100 types of QW dynamics, and the histograms of the infidelity and purity error of the reconstructed quantum states are displayed in Figs. 7(b) and 7(c), respectively. The corresponding average values of the infidelity and purity error are $0.025\pm 0.012$ and $0.042\pm 0.037$, respectively, indicating that the experimental states align well with the theoretical expectations.

### v3

@@ Abstract
The growing maturity of photonic integrated circuit (PIC) fabrication technology enables the high integration of an increasing number of optical components onto a single chip. With the incremental circuit complexity, the calibration of active phase shifters in a large-scale PIC becomes a crucially important issue. The traditional one-by-one calibration techniques encounter significant hurdles with the propagation of calibration errors, and achieving the decoupling of all phase shifters for independent calibration is not straightforward. To address this issue, we propose a global calibration approach for large-scale PIC. Our method utilizes a custom network to simultaneously learn the nonlinear phase-current relations for all thermo-optic phase shifters on the PIC by minimizing the negative likelihood of the measurement datasets. Moreover, the reflectivities of all static beams plitter components can also be synchronizedly extracted using this calibration method. As an example, a quantum walk PIC with a circuit depth of 12 is calibrated, and a programmable discrete-time quantum walk is experimentally demonstrated. These results will greatly benefit the applications of large-scale PICs in photonic quantum information processing.
@@ I Introduction
The traditional calibration method requires the decoupling of all phase shifters followed by individual calibration for each one [19, 20, 21, 22, 23]. However, this approach is not ideal for large-scale PICs due to two main challenges. Firstly, while the traditional method can be used to calibrate the standard triangular “Rech” PICs [24, 25], rectangular “Clements” PICs [26, 27, 28, 29], and PICs with cascaded phase shifters [30, 31] by following a specific decoupling order, it is difficult to find an effective and direct way to decouple all phase shifters in the intricate mesh structures. Secondly, deviations in the calibration values of phase shifters in preceding levels may also spread errors in the calibration of subsequent phase shifters. As a result, calibration errors accumulate progressively with the increasing number of phase shifters, eventually becoming significant and non-negligible.
To address these issues, we present an efficient calibration method developed to globally calibrate all optical components on the PICs, thereby mitigating the accumulation of errors inherent in the traditional one-by-one calibration approach. The global calibration method utilizes a custom network with simulation layers that creates a virtual duplicate of the physical PICs, modelling their practical nonunitary processes involving optical loss. Using the gradient-descent approach, we can determine the optimal values of parameters that need to be calibrated, ensuring that the outputs of the network closely resemble the distributions of measurement data. As an example, on a large-scale quantum walk (QW) PIC with a circuit depth of 12, we find that the trained custom network enables the simultaneous extraction of not only the nonlinear phase-current relations of all the thermo-optic phase shifters but also the reflectivities of all the beam splitters. Moreover, we further experimentally demonstrate the silicon PIC calibration and implement the on-chip QW [32, 33, 34, 35, 36, 37, 38, 39, 40].
This paper is organized as follows: In Sec. II, we will introduce our custom-architecture global calibration method, provide a concise overview of the QW PIC, and evaluate the performance of our calibration method using synthetic datasets featuring Gaussian noise. In Sec. III, we will showcase the calibration results of a silicon PIC utilizing noisy experimental datasets and discuss the performance regarding the experimental implementation of QW. Finally, Sec. IV presents the summary of our work and engages in a discussion.
@@ II.1 Global calibration method
where $\sin^{2}\theta$ denotes the reflectivity of the beam splitter. Considering the inevitable optical loss in a practical chip, the transfer matrix for a lossy BS becomes nonunitary and is given by [41]:
@@
where $\alpha$ denotes the transfer efficiency of a single BS, and $1-\alpha$ represents its optical loss. The phase shifter provides a relative phase difference $\phi$ between the two output paths of the beam splitter, and its matrix representation is $\hat{U}_{\text{PS}}=\text{diag}\{1,e^{i\phi}\}$. For the widely used thermo-optic phase shifter, the relative phase $\phi$ can be tuned by the current $I$ applied to the phase shifter, according to the following formula [13]:
@@
where $a$ and $b$ are hardware parameters to be determined. Then, the nonunitary process $\hat{U}(\bm{a},\bm{b},\bm{\theta},\bm{\alpha})$ of PICs is the matrix product of its individual components. Moreover, considering the mode-field diameter mismatch between the output ports of the PICs and the optical fiber arrays, it is necessary to introduce another parameter, $\bm{\eta}$, to indicate the fiber-chip coupling efficiency. Consequently, $\bm{\lambda}=\{\bm{a},\bm{b},\bm{\theta},\bm{\alpha},\bm{\eta}\}$ are all undetermined parameters for a linear photonic circuit.
@@
The learning process necessitates a number of training data, comprising randomly generated current distributions $\{\bm{I}\}_{n}$ applied to all the phase shifters and the target probability distributions $\{\bm{P}_{\text{tar}}\}_{n}$ of the output ports on the chip. The index $n=1,\cdots,N$ represents each training sample, and $N$ is the total number of training datasets. We set the current distribution $\bm{I}$ as inputs, which are then fed to the input layer of our custom network. The custom-network-parameterized nonunitary process $\hat{U}_{\text{ML}}(\bm{a},\bm{b},\bm{\theta},\bm{\alpha})$ can be determined using Eqs. (2) and (3). For a given initial state $\rho_{i}$, the final state is given by $\rho_{f}=\hat{U}_{\text{ML}}\rho_{i}\hat{U}_{\text{ML}}^{\dagger}$, and the probability distribution $\text{Tr}[\left|v\right\rangle\!\!\left\langle v\right|\rho_{f}]$ can be obtained. Considering the relative fiber-coupling efficiency $\bm{\eta}$ between the output ports of the PICs, the final probability distribution is normalized, and then the parameterized network distribution $\bm{P}_{\text{ML}}(\bm{\lambda})$ is generated.
We define the cost function $\mathcal{L}$ as the summation of the L1-norm distance [42] between the target probability distributions in the training datasets $\{\bm{P}{{}_{\text{tar}}}\}_{n}$ and the parameterized network distributions $\{\bm{P_{\text{ML}}(\bm{\lambda})}\}_{n}$: $\mathcal{L}(\bm{\lambda})=\sum_{n}\frac{1}{2}|\bm{P{{}_{\text{ML}}}}(\bm{\lambda})-\bm{P_{\text{tar}}}|$. The aim of training the custom network is to find the optimal parameters $\bm{\lambda}^{\ast}$ by minimizing the cost function such that the parameterized network distributions approximate the target ones. Then, the parameterized nonunitary process $\hat{U}_{\text{ML}}$ together with the relative coupling efficiency $\bm{\eta}^{\ast}$ for the trained network can well reflect the target physical process that generates the distributions $\{\bm{P_{\text{ML}}}(\bm{\lambda}^{\ast})\}_{n}\simeq\{\bm{P}{{}_{\text{tar}}}\}_{n}$.
@@ II.2 Model of on-chip quantum walks
Here, the parameter $\alpha$ determines the loss strength and varies depending on both the position and time step. $\cos^{2}\theta$ and $\sin^{2}\theta$ denote the probabilities of a walker moving backwards and forwards on the 1D line, respectively. Parameters $\beta$ and $\gamma$ introduce the relative phase difference. The coin operator $\hat{C}$ can be position and time dependent, a feature necessary for specific tasks in QW-based quantum simulations [35] and quantum algorithms [14].
For the QW PIC, the coin state {$\{\ket{\uparrow},\ket{\downarrow}\}$ in the QW model can be cast on the path state (inset of Fig. 2): the upper path of the beam splitter corresponds to |$\ket{u}$, while the lower path corresponds to |$\ket{l}$. The localized coin state at each position is described by |$\ket{c}=c_{1}\ket{u}+c_{2}\ket{l}$, which can be interpreted as a superposition state located on the two paths of the corresponding beam splitter. Each beam splitter with a splitting ratio of 50:50 and phase shifter work together (black circle in Fig. 2) to realize the coin operator $\hat{C}$ in the sense of path states, corresponding to Eq. (4) with $\theta=\pi/4$ and $\beta=0$. The interwoven mesh architecture of the PIC, composed of the on-chip waveguides (black solid lines), realizes the shift operators $\hat{S}$. After $T$-step nonunitary dynamics of the QW PIC, the size of the position space is $T$, with each position having two paths representing the coin space. Therefore, the total number of output ports of the QW PIC is $2T$.
@@ II.3 Numerical benchmarks
During the iteration process, simultaneously updating all network parameters using the gradient-descent method is susceptible to getting trapped in a local minimum of $\mathcal{L}$. To mitigate this issue and escape local minima, we utilize a nonsimultaneous parameter update method, updating only partial parameters at each epoch while keeping the remaining parameters fixed. The relative coupling efficiency $\bm{\eta}$ is configured to be constantly updated. The reflectivity parameter $\bm{\theta}$ and the relative transfer efficiency $\bm{\alpha}$ are frozen at first. We then alternatively train one of the two parameters $\{\bm{a},\bm{b}\}$ while the remaining one is kept frozen and untrainable. Within a single alternating training process, the training for each parameter ceases upon meeting the cut-off condition, where five consecutive decreases in the loss function between adjacent epochs are each less than $10^{-5}$. After eight rounds of alternating training, we proceed to train the last two parameters, $\bm{\theta}$ and $\bm{\alpha}$, while maintaining the fixed values of $\bm{a}$ and $\bm{b}$. In the end, following the completion of the nonsimultaneous parameter updates described above, we proceed to simultaneously update all parameters $\bm{\lambda}$ until the cut-off condition is met.
@@ III.1 Experimental setup
In our experiment, the conceptual diagram of the QW PIC depicted in Fig. 2 is implemented through silicon waveguide circuits fabricated using the CMOS-compatible silicon photonics process [51], as illustrated in Fig. 5. The $6.8\,\text{mm}\times 3.4\,\text{mm}$ chip comprises an array of 66 individually controlled thermo-optic phase shifters with 78 waveguide beam splitters to form the cascade of the interference network of a 12-step on-chip QW. The controlled phase shifter is implemented by electrically heating a resistor to change the refractive index of the waveguide to deform the optical path length [52]. Then, the phase shifter has a linear phase-power relation, and its phase change $[0,2\pi]$ can be tuned by changing the imposed current according to the nonlinear phase-current relation in Eq. (3). To program all these phase shifters, we have developed a multichannel electronic control system. The beam splitter elements are realized using multimode interference couplers. On-chip grating couplers [53] are used for fiber-chip coupling, and the coupling loss is 4.5 dB per facet. A continuous-wave laser (TLX1, Thorlabs) operating at a wavelength of 1550 nm and with a power output of 5 mW is coupled into the QW PIC through a single-mode optical fiber. The input fiber is precisely aligned with the lower path of the first waveguide beam splitter using a three-axis motorized positioning stage (KWL06050, Suruga), and the polarization of the input laser is fine-tuned using the three-paddle fiber polarization controller (FPC030, Thorlabs). We simultaneously measure the light intensity from all output ports by coupling them to an array of optical fibers and subsequently directing them to Germanium-amplified detectors (PDA30B2, Thorlabs). For each output port, we detect the light intensity five times and store the average value. Consequently, the measured intensity distributions are normalized to the experimental probability distribution $\bm{P}_{\text{exp}}$. In the experiment, the QW PIC is temperature controlled at 22C.
@@ III.2 Experimental results
Herein, the calibration of the QW PIC is implemented and tested by training the custom network directly on actual experimental datasets $\{\bm{P}_{\text{exp}}\}_{n}$. In order to generate the experimental training and test datasets, we randomly generate 1500 sets of current distributions $\{\bm{I}\}_{n}$ ranging from 0 mA to 7 mA imposed on the phase shifters on our silicon PIC. We obtain 1500 sets of experimental probability distributions $\{\bm{P}_{\text{exp}}\}_{n}$ after normalizing the measured light intensity over all output ports. The experimental datasets are divided into two parts according to an 85:15 ratio, where 1275 sets of datasets are used to train the custom network, and the remaining 225 sets of datasets are used to test the trained network. With these experimental training sets, the parameters $\bm{\lambda}=\{\bm{a},\bm{b},\bm{\theta},\bm{\alpha},\bm{\eta}\}$ of the custom network in Fig. 1 are then pretrained.
The trained network can characterize the phase-current relation, assumed to be of the form $\phi=aI^{2}+b$, the reflectivity parameter $\bm{\theta}$ and the relative transfer efficiency $\bm{\alpha}$ of BSs, as well as the relative output coupling efficiency $\bm{\eta}$ between the output ports. Here, we provide the relative transfer efficiency between all BSs and the relative coupling efficiency between all output ports because we utilize the normalized probability distributions to calibrate the QW PIC. Even so, we need only to measure the absolute transfer efficiency for one BS and the absolute coupling efficiency for one output port. The efficiencies of all the remaining BSs and output ports can then be directly calculated.
@@
Figure 7: Validation of the QW PIC. (a) The measured probability distribution $\bm{P}^{\text{QW}}_{\text{exp}}$ (blue bars) for 12-step Hadamard QWs. The red bars mean the distribution $\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})$ predicted by the trained network. (b) and (c) display histograms of the measured values of infidelity and purity error for 100 output quantum states after nine-step QW dynamics, respectively.
@@
After completing the calibration of the photonic chip using the training datasets, we utilize the experimental test datasets to further verify the accuracy of the calibration. We apply the remaining 225 sets of distributions of random currents, and compare the experimentally measured distributions $\{\bm{P}^{\text{test}}_{\text{exp}}\}_{n}$ with the predicted probability distributions $\{\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})\}_{n}$ of the trained network with the optimal hyperparameters $\bm{\lambda}^{\ast}$. The difference between the two distributions for each test sample $n$ is defined by L1-norm distance: $d(\bm{\lambda}^{\ast})=\frac{1}{2}|\bm{P{{}_{\text{ML}}}}(\bm{\lambda}^{\ast})-\bm{P}^{\text{test}}_{\text{exp}}|$. We observe that the distance values for $85\%$ of the test samples are less than 0.05, as shown in Fig. 6(f). The average distance for all test samples is $\bar{d}=0.033\pm 0.018$, indicating that our calibration results for the phase shifters and beam splitters are in good agreement with the actual physical process. The slight deviation suggests the presence of additional chip imperfections that are not considered by our custom network, such as the nondominant thermal crosstalk effect [54]. The error bar of $\bar{d}$ represents one standard deviation.
@@
After verifying the performance on numerous random validation samples using the calibrated photonic chip, we proceed to implement the conventional 1D Hadamard QWs, where the coin operators are fixed as the Hadamard gate and independent of time and position. We inject 1550-nm light into the lower path of the first waveguide beam splitter followed by setting the phase of the first phase shifter to $\pi/2$. Therefore, we prepare a balanced superposition initial state between two paths with a phase difference of $\pi/2$ through one-step QW dynamics [55]. The phases of the remaining phase shifters are set to 0. We measure the light intensity coupled to the optical fiber arrays, and the experimental probability distribution $\bm{P}^{\text{QW}}_{\text{exp}}$ for a 12-step QW is obtained using the calibrated fiber-coupling efficiency for compensation, as depicted in Fig. 7(a). The resulting L1-norm distance with respect to the QW distribution predicted by the trained network is $d=0.057$, indicating their good agreement.
@@
Beyond the position distributions, we then perform the quantum state tomography for the calibrated QW PIC with inherent high-dimensional structures [56, 57, 58, 59]. To achieve this, in addition to the computational basis |$\ket{v}$, extra measurement bases need to be introduced to fully characterize the output quantum state. We divide the QW PIC into two sections: the first part, with a circuit depth of nine, is used to implement arbitrary nine-step QW dynamics, while the second part, with a circuit depth of 3, serves as a high-dimensional measurement device [60, 61, 62, 63]. For each nine-step QW dynamics, we impose 14 sets of current distributions into the measurement device and measure the output probability distributions. Here, we perform standard maximum-likelihood reconstruction with the estimator taking the form ${\rho_{\text{m}}=TT^{\dagger}/\text{Tr}(TT^{\dagger})}$, where $T$ is a lower triangular matrix [64]. The quantum states after the 9-step QW can be reconstructed by maximizing the likelihood of 14 sets of experimental distributions. We arbitrarily generate 100 types of QW dynamics, and the histograms of the infidelity and purity error of the reconstructed quantum states are displayed in Figs. 7(b) and 7(c), respectively. The corresponding average values of the infidelity and purity error are $0.025\pm 0.012$ and $0.042\pm 0.037$, respectively, indicating that the experimental states align well with the theoretical expectations.